# On Geometry-Enhanced Parameter-Efficient Fine-Tuning for 3D Scene Segmentation

### v1

@@ 5 Conclusion
While GEM achieve promising performance with little overhead in complexity, we notice it relies on the assumptaion that the model can be fully parallelized. For example, it would incur noticeable overhead if using large batch sizes, as the device is forced to process the input sequentially, in sipte of the parallel design in GEM. In addition, we are able to successfully train our PEFT methods with a fractional of epoch than the full fine-tining, e.g. 100 ranther than 800 epochs, we realize that the gradients need to be back-propagated into early layer and could hinder the training efficiency, especially with dense point cloud like S3DIS [1]. We thus suggest that it is required a systematic examination on the additional overhead that PEFT methods would experience in 3D scene understanding.
@@
In this supplementary, we provide more materials as follow,
Appendix A  details the implementation, training, and inference;
Appendix B  presents additional experiments and analyses with tight parameter budgets
Appendix C  offers more qualitative visualizations.
Appendix A Details of Implementation, Training and Inference
Our implementation is based on the open-source codebase Pointcept (here) and follows the official implementations for Sonata [74], PPT [77] as well as PTv3 [75].
Leveraging the parameter efficiency of our method, we train on a single 4090 GPU for much fewer epochs to obtain the reported performance. For example, we train on ScanNet for 100 epochs, in contrast to the 800 epochs in the released Sonata configuration. Our code will also be released.
Table LABEL:tbl:profiling summarizes empirical latency and memory usage. We intentionally refrain from deployment-specific optimizations such as LoRA weight merging. Although the global attention and local convolution modules introduce extra cost, the runtime and memory footprint stay within the same order of magnitude as other PEFT baselines.
@@ Appendix B More Experiments and Analysis
In spite of the promising results shown in Sec. 3, we suggest that it can better demonstrates the full potentials of PEFT methods under more challenging settings.
To stress parameter efficiency, we constrain all methods to the same limited learnable-parameter budgets, including enforcing rank $r=1$, allowing 0.1% parameters, and 1% parameters. As reported in LABEL:tbl:semseg_tight, GEM delivers consistently higher accuracy, whereas several baselines fail when the budget becomes extremely stringent.
@@ Appendix C More Visualizations
We provide more qualitative results regarding the attention map generated by our latent tokens in LABEL:fig:attn-more. Despite shared across the stages, the latent tokens appear to focus on differen parts of the scenes, providing different scene context to the backbone models at different stages.

### v2

@@ 5 Conclusion
While GEM achieves promising performance with little overhead in complexity, we notice it relies on the assumption that the model can be fully parallelized. For example, it would incur noticeable overhead if using large batch sizes, as the device is forced to process the input sequentially, in spite of the parallel design in GEM. In addition, we are able to successfully train our PEFT methods with a fraction of epochs than the full fine-tuning, e.g. 100 rather than 800 epochs, we realize that the gradients need to be back-propagated into early layers and could hinder the training efficiency, especially with dense point clouds like S3DIS [s3dis]. We thus suggest that it is required a systematic examination on the additional overhead that PEFT methods would experience in 3D scene understanding.
Acknowledgement.
Dr. Tao’s research is partially supported by NTU RSR and Start Up Grants.
@@ Acknowledgement.
In this supplementary, we provide more materials as follows,
Appendix˜A details the implementation, training, and inference;
Appendix˜B presents additional experiments and analyses with tight parameter budgets
Appendix˜C offers more qualitative visualizations.
Appendix A Details of Implementation, Training, and Inference
Our implementation is based on the open-source codebase Pointcept (here) and follows the official implementations for Sonata [sonata], PPT [ppt], as well as PTv3 [ptv3].
Leveraging the parameter efficiency of our method, we train on a single 4090 GPU for much fewer epochs to obtain the reported performance. For example, we train on ScanNet for 100 epochs, in contrast to the 800 epochs in the released Sonata configuration. Our code is available at https://github.com/LiyaoTang/GEM.
Tab.˜8 summarizes empirical latency and memory usage. We intentionally refrain from deployment-specific optimizations such as LoRA weight merging. Although the global attention and local convolution modules introduce extra cost, the runtime and memory footprint stay within the same order of magnitude as other PEFT baselines.
@@ Appendix B More Experiments and Analysis
In spite of the promising results shown in Sec.˜3, we suggest that it can better demonstrate the full potential of PEFT methods under broader and more challenging settings.
PEFT for shape analysis.
While GEM is motivated by PEFT for large-scale 3D scenes, most existing 3D PEFT methods target object-level understanding, as discussed in Sec.˜2. To assess cross-domain generalization and compare against these existing 3D-specific PEFT methods, we further evaluate GEM on 3D shape datasets, ShapeNetPart [shapenetpart]. As shown in Tab.˜8, GEM remains competitive and achieves the best performance. We also observe that common backbones for shape analysis, such as ReCon [shape_recon], attach large MLP heads that account for approximately 20% of all parameters, rendering fine-tuning inefficient. To broaden the applicability of backbone models for 3D shapes, an interesting direction for future work is to develop architectures with more parameter-efficient heads that are better suited to fine-tuning.
PEFT for 3D convolutional networks.
While our primary experiments apply GEM to state-of-the-art backbones, mainly transformer models, we also verify its generality on convolutional architectures. In particular, we integrate GEM into SparseUNet [minkowski], a 3D convolutional model, following the MSC [msc] protocol to pre-train on ScanNet [scannet] and fine-tune on ScanNet200 [scannet200]. As reported in Tab.˜10, GEM improves performance by +4.6 mIoU over linear probing and outperforms other PEFT baselines, demonstrating robustness beyond transformer models. We notice the gains are narrower than those on transformer backbones, largely due to the limited capacity of convolutions, as also indicated by a near-collapse linear probing performance in this setting.
PEFT under tight budgets.
To stress parameter efficiency, we constrain all methods to the same limited learnable-parameter budgets, including enforcing rank $r=1$, allowing 0.1% parameters, and 1% parameters. As reported in Tab.˜10, GEM delivers consistently higher accuracy, whereas several baselines fail when the budget becomes extremely stringent.
@@ Appendix C More Visualizations
We provide more qualitative results regarding the attention map generated by our latent tokens in Fig.˜4. Despite shared across the stages, the latent tokens appear to focus on different parts of the scenes, providing different scene context to the backbone models at different stages.
Methods Params. Memory Latency
Base model
Sonata (full.) [sonata] 124.8M 8.2G 61.8ms
Sonata (ft.) 108.5M 6.4G 50.2ms
PEFT methods
Sonata (lin.) 0.02M 6.4G 50.2ms
+ Adapter [adap] 2.8M 6.7G 50.6ms
+ LoRA [lora] 1.9M 7.9G 52.8ms
+ Prompt Tunning [prefix] 5.5M 7.3G 51.5ms
+ GEM (ours) 1.8M 7.1G 57.8ms
Table 7: Inference efficiency of PEFT methods. We benchmark with the batch size fixed to 1.
Shape-Part Seg. Params ShapeNetPart [shapenetpart]
Methods Learn. Cls. mIoU Inst. mIoU
Full fine-tuning
ReCon (ft.) [shape_recon] 27.06 84.52 86.1
PEFT methods
ReCon (lin.) [shape_recon] 5.23M 83.06 85.2
+ PointLoRA [ptpeft_pointlora] 5.63M 83.98 85.4
+ PointGST [ptpeft_gst] 5.59M 83.98 85.8
+ GEM (ours) 5.58M 84.02 85.8
Table 8: Generalizing to 3D shapes.
Figure 4: Attention maps of our latent tokens in context adapter.
Tight Budgets Params ScanNet200 Val [scannet200]
Methods Learn. Pct. mIoU mAcc allAcc
Training from scratch
SparseUnet [minkowski] 39.2M 100% 25.0 32.9 80.4
Full fine-tuning
SparseUnet (ft.) [msc] 0.02M 0.05% 32.0 41.6 82.3
PEFT methods with rank=1
SparseUnet (lin.) 0.02M 0.05% 1.5 2.5 53.6
+ BitFit [select_bitfit] 0.03M 0.07% 4.8 6.9 62.7
+ Adapter [adap] 0.6M 1.5% 9.5 13.4 69.5
+ LoRA [lora] 0.8M 2.0% 13.2 17.6 74.7
+ GEM (ours) 0.6M 1.5% 15.2 22.9 75.6
Table 9: Generalizing to convolutional networks.
Tight Budgets Params ScanNet Val [scannet]
Methods Learn. Pct. mIoU mAcc allAcc
Training from scratch
PTv3 [ptv3] 124.8M 100% 77.6 85.0 92.0
Full fine-tuning
Sonata (full.) [sonata] 124.8M 100% 79.4 86.1 92.5
Sonata (ft.) 108.5M 100% 78.3 85.9 92.3
PEFT methods with rank=1
Sonata (lin.) 0.02M 0.02% 72.5 83.1 89.7
+ Adapter [adap] 0.05M 0.04% 74.0 84.1 90.5
+ LoRA [lora] 0.05M 0.05% 42.5 55.4 75.1
+ Prompt Tunning [prefix] 0.05M 0.05% 72.9 83.5 90.0
+ GEM (ours) 0.07M 0.06% 75.2 84.8 91.1
PEFT methods with 0.1% params.
Sonata (lin.)
+ Adapter [adap] 0.1M 0.1% 75.1 84.7 91.1
+ LoRA [lora] 0.1M 0.1% 75.0 84.4 91.1
+ Prompt Tunning [prefix] 0.1M 0.1% 73.5 83.9 90.3
+ GEM (ours) 0.1M 0.1% 76.5 85.5 91.6
PEFT methods with 1% params.
Sonata (lin.)
+ Adapter [adap] 1.1M 1.0% 76.6 85.3 91.7
+ LoRA [lora] 1.1M 1.0% 76.7 85.6 91.7
+ Prompt Tunning [prefix] 1.1M 1.0% 73.8 84.2 90.4
+ GEM (ours) 1.1M 1.0% 78.2 86.3 92.2
Table 10: Performance with tight budgets.